OpenStudy (anonymous):

how to do: 1/2x-5+2x+3=10

OpenStudy (anonymous):

4.8

OpenStudy (anonymous):

1/2x-5+2x+3=10 =>[1-5(2x)+2x(2x)+3(2x)]/2x=10 =>(1-10x+4x^2+6x)/2x=10 =>4x^2-4x+1=20x =>4x^2-24x+1=0 a=4 b=-24 c=1 So, \[[24\pm \sqrt{(-24)^{2}-4(4)1}]/2(4)\] As we can see the Discriminant is 23.66. So x is either 5.9575 or 0.0425